# ControlVAR: Exploring Controllable Visual Autoregressive Modeling

### v1

@@ Abstract
Conditional visual generation has witnessed remarkable progress with the advent of diffusion models (DMs), especially in tasks like control-to-image generation. However, challenges such as expensive computational cost, high inference latency, and difficulties of integration with large language models (LLMs) have necessitated exploring alternatives to DMs. This paper introduces ControlVAR, a novel framework that explores pixel-level controls in visual autoregressive (VAR) modeling for flexible and efficient conditional generation. In contrast to traditional conditional models that learn the conditional distribution, ControlVAR jointly models the distribution of image and pixel-level conditions during training and imposes conditional controls during testing. To enhance the joint modeling, we adopt the next-scale AR prediction paradigm and unify control and image representations. A teacher-forcing guidance strategy is proposed to further facilitate controllable generation with joint modeling. Extensive experiments demonstrate the superior efficacy and flexibility of ControlVAR across various conditional generation tasks against popular conditional DMs, e.g., ControlNet and T2I-Adaptor.
@@ 1 Introduction
In recent years, conditional image generation [1, 2, 3, 4, 5] has attracted great attention and there have been significant advancements in text-to-image generation [6, 7, 8], image-to-image generation [1, 2, 9], and even more complex tasks [5, 10, 11]. Most recent approaches, e.g., ControlNet [1], leverage the powerful diffusion models (DMs) [6, 12] to model the large-scale image distribution and incorporate additional controls with classifier-free guidance [13]. However, the inherent nature of the diffusion process imposes many challenges for the diffusion-based visual generation: (1) the computational cost and inference time are significant due to the iterative diffusion steps [14, 15] and (2) the incorporation in mainstream intelligent systems, i.e., large language models (LLMs) [16, 17], is intricate due to the representation difference. This motivates the community to find a replacement for DMs for high-quality and efficient visual generation in the era of LLMs.
@@
Inspired by the success of autoregressive (AR) language modeling [16, 17], AR visual modeling [3, 4] has been investigated as a counterpart to DMs given its strong scalability and generalizability [4, 18]. Several inspiring works, e.g., VQGAN [3], DALL-E [19] and VAR [4], have demonstrated promising image generation results with AR modeling. Nevertheless, compared to the prosperity of conditional DMs [1, 2, 20, 21, 22, 23], visual generation with conditional AR modeling [24, 3] remains significantly under-explored. Different from DMs, where all the pixels are modeled simultaneously, AR models are characterized by modeling sequential values based on their corresponding previous ones. This AR approach naturally leads to a conditional model, providing potential flexibility when incorporating additional controls. To leverage this property, teacher forcing is a popular approach that controls AR prediction by replacing partially predicted tokens with ground truth ones [3]. Thanks to this nature of AR modeling, we found that highly flexible conditional generation can be achieved by teacher forcing partial AR sequence with proper model designs.
@@
Figure 2: In contrast to previous methods [3, 24] that leverage prefix conditional tokens to impose controls, ControlVAR jointly models the pixel-level controls and image during training and conducts the conditional generation tasks during testing with the teacher forcing. Class and type tokens provide semantic and control type (mask, canny, depth and normal) information respectively.
@@
In this paper, we explore the Controllable Visual AutoregRessive modeling with both token-level and pixel-level conditions. A new conditional AR paradigm, ControlVAR is introduced, which permits a highly flexible conditional image generation by embracing the next-scale prediction of joint control and image (Fig. 2(b)). Previous wisdom [24, 3] typically utilizes prefix conditions (Fig. 2(a)) and mainly model images from raw pixel space in an AR manner. Differently, we notice that if we jointly model the control and image, the learned joint prediction can be easily guided by teacher forcing during inference. On the one hand, we unify the control and image representations and reformulate the sequential variables for the AR process to enable effective joint modeling. On the other hand, by analyzing the modeled probabilities, we introduce an effective sampling strategy, named teacher forcing guidance (TFG) to facilitate conditional sampling. Remarkably, a single ControlVAR model trained via TFG is capable of multiple meaningful tasks with different input-output combinations between control and image: (a) joint control-image generation, (b) control/image completion, (c) control-to-image generation, (d) image-to-control generation, as demonstrated in Fig. 1. Beyond the image-control tasks that are jointly modeled during training, we observe that ControlVAR also emerges capabilities for unseen tasks, e.g., control-to-control generation, further enhancing its flexibility and versatility. Our contribution can be summarized in three-fold:
@@ 2.1 Diffusion-based Image Generation
The evolution of diffusion models, initially introduced by Sohl-Dickstein et al. [25] and later expanded into image generation using fixed Gaussian noise diffusion processes [15, 26], has witnessed significant advancements driven by various research efforts. Nichol et al. [27] and Dhariwal et al. [28] proposed techniques to enhance the effectiveness and efficiency of diffusion models, paving the way for improved image generation capabilities. Notably, the paradigm shift towards modeling the diffusion process in the latent space of pre-trained image encoders as a strong prior [29, 3] rather than raw pixels spaces [30, 31, 12] has been instrumental in achieving high-quality image generation with reasonable inference speed. This approach has led to the development of foundational diffusion models such as Glide [32], Cogview [33, 34, 35], Make-a-scene [36], Imagen [37], DALL.E [38], Stable Diffusion [39], MidJourney [40], SORA [41], among others, which are often pre-trained on large-scale data with conditions, typically text [42, 43, 44, 45]. Recent advancements include consistency models derived from diffusion models [46, 47, 48], enabling generation with reduced inference steps. These foundational diffusion models have not only opened doors to novel downstream applications like Text inversion [8], DreamBooth [9], T2I-Adapter [2], ControlNet [1], but also inspired a plethora of research in controllable generation [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 24] and other innovative areas.
@@
Unlike diffusion-based models that typically leverage continuous image representation, autoregressive models [60, 3, 61, 4] leverage discrete image tokens. An image tokenizer [3, 62, 63, 64, 65] is utilized to encode the image into a sequence of discrete tokens. VQGAN [3] first patches the image and then employs a vector-quantization approach to discretize the image features. Following this paradigm, a series of following-up works improve the image tokenization by using more powerful quantization operations [60, 66, 62], reformulating the image representation [4, 67] and modifying the network architecture [68, 69]. With the discrete tokens, a transformer structure [70] is leveraged to model the image token sequences. RQ-GAN [66] improves the modeling by incorporating a hierarchy design and MQ-VAE [60] further utilizes StackTransformer to enhance the spatial focus. MUSE [7] is a large-scale pre-trained text-to-image model where a low-resolution image is first generated followed by a super-resolution transformer to refine the image. Recently, VAR [4] introduced a new next-scale autoregressive prediction paradigm where the image representation is shifted from patch to scales. The new representation is featured with the maintenance of spatial locality and much lower computational cost. In this paper, we follow the next-scale autoregressive paradigm and explore the incorporation of additional controls into the modeling process.
@@ 2.3 Conditional Image Generation
Though significant progress has been made in generating highly realistic images from textual descriptions, describing every intricate detail of an image solely through text poses challenges. To overcome this limitation, researchers have explored alternative approaches using various additional inputs to effectively control image and video diffusion models. These inputs encompass bounding boxes [58, 71], reference object images [9, 72], segmentation maps [73, 74, 1], sketches [1], and combinations thereof [75, 22, 76, 77, 78, 5, 79]. However, fine-tuning the vast array of parameters in these diffusion models can be computationally intensive. To address this, methods like ControlNet [1] have emerged, enabling conditional control through parameter-efficient training strategies [1, 80, 2]. Notably, X-Adapter [81] innovatively learns an adapter module to adapt ControlNets pre-trained on smaller image diffusion models (e.g., SDv1.5) for larger models (e.g., SDXL). SparseCtrl [82] takes a different approach, guiding video diffusion models with sparse conditional inputs, such as few frames instead of full frames, to mitigate the data collection costs associated with video conditions. However, the implementation of SparseCtrl necessitates training a new variant of ControlNet from scratch, as it involves augmenting ControlNet with an additional channel for frame masks. Beyond traditional conditional image generation, the in-context learning capability of conditional models has also been explored [83, 78, 18, 84]. LVM [18] investigates the scaling learning capability of a large vision model without any linguistic data. 4M [78] investigate the large-scale visual generation with multimodal data using masked image modeling. Different from previous works which are mainly focusing on diffusion models, we aim to explore adding additional control to the autoregressive visual generation process.
@@ 3 ControlVAR
ControlVAR is an autoregressive Transformer [85] framework for conditional image generation tasks, using the following as conditions: image $I\in\mathbb{R}^{3\times H\times W}$, pixel-level control $C\in\mathbb{R}^{3\times H\times W}$ and token-level control $c\in\mathbb{R}^{D}$ where $H,W$ and $D$ denotes the image size and dimension of control token respectively. We denote the set of $N$ different types of controls as $\mathcal{C}=\{C_{n}\}_{n\in[N]}$.
@@
Prior conditional approaches [1, 4] have often utilized distinct models for individual control type $C$, learning a conditional distribution in the form of $p(I|C,c)$, where each image $I$ is encoded as a sequence of discrete tokens of length $T$, denoted as $(x_{1},x_{2},\dots,x_{T})$. By employing autoregressive (AR) modeling, we can rewrite the conditional probability $p(I|C,c)$ as
@@
We consider four popular control types - entity mask, canny, depth, and normal in this paper. We notice that canny, depth, and normal can be easily converted to RGB by using simple transformations [1]. However, entity segmentation masks $M\in\{0,1\}^{N\times H\times W}$ which comprises $N$ class-agnostic binary masks (Fig. 3(b)) cannot be easily converted.
@@
Inspired by SOLO [86], we leverage a position-aware color map to encode the binary masks $M$ into a colormap $M^{\prime}\in[0,255]^{3\times H\times W}$. To better distinguish the color difference, we select 5 candidate values $\{0,64,128,192,255\}$ from each RGB channel and combine them to $1124=5^{3}-$ 1 colors ($(0,0,0)$ is preserved for background). To apply the colormap, as shown in Fig. 3, we divide the image into $n_{h}\times n_{w}$ regions where each region represents a corresponding color. We calculate the centeredness of each mask and apply the colors to masks based on their centeredness locations. Therein, we can convert the entity masks to a RGB colormap.
@@
As the control and image share the same RGB representation, we can utilize the same approach to tokenize them. To represent an RGB image as a sequence of discrete tokens $(x_{1},x_{2},\dots,x_{T})$, patch-level [3] and scale-level [4] representations have been explored. The patch-level tokenization process splits an image into $T$ patches and represents each patch as a token $x_{t}$ where $x_{t}\in[V]^{1}$ is an integer from a vocabulary of size $V$. Recently, a scale-level representation has been introduced which decomposes the image into $T$ scales where each scale is represented by a set of tokens $x_{t}\in[V]^{h_{t}\times w_{t}}$ (Fig. 2(b)). $h_{t}\times w_{t}$ denotes the size of the $t$-th scale. Compared to patch-level representation, scale-level representation can better preserve the spatial locality and capture global information which are desired for conditional image generation tasks. This motivates us to adopt the scale-level representation in our approach. Specifically, we obtain the image tokens and control tokens using the shared tokenizer $\Phi$ as
@@ 3.2 Joint Control-Image Modeling
We demonstrate the network details for joint modeling in this section. Following VAR [4], we leverage a GPT-2 style Transformer network architecture for our ControlVAR models. As shown in Fig. 4 (a), we jointly model the control and image in each stage. A flatten operation is adopted to convert the sequence of 2D features into 1D. Full attention is enabled for both control and image tokens belonging to the same scale, which allows the model to maintain spatial locality and to exploit the global context between control and image. A standard cross entropy loss is used to supervise our autoregressive ControlVAR models.
Specifically, we employ two pre-defined special tokens $c=\texttt{[CLS]}\in[N_{cls}]^{1}$ and $c_{t}=\texttt{[TYP]}\in[N_{typ}]^{1}$ as the start tokens. $N_{cls}$ and $N_{tpy}$ denote the number of classes and control types respectively. [CLS] token aims to provide semantic context for the generated image. [TYP] token is used to select the type of pixel-level control to be generated along with the image. Following previous works [7, 4], additional empty tokens are used to replace special tokens with a probability of $\delta$ during training to apply classifier-free guidance [13].
@@
Classifier-free guidance (CFG) [13] was originally introduced to apply and enhance the effect of conditional controls on diffusion models without an explicit classifier. Extensive studies [87, 7, 4] have demonstrated that classifier-free guidance also works for AR models.
@@
Here, we analyze how to achieve conditional generation by using the image generation task $p(I|C,c,c_{t})$ as an example. Given image $I$, pixel-level control $C$ and token-level controls $c,c_{t}$, CFG [13] leverages Bayesian rule to rewrite the conditional distribution as

### v2

@@ Abstract
Conditional visual generation has witnessed remarkable progress with the advent of diffusion models (DMs), especially in tasks like control-to-image generation. However, challenges such as expensive computational cost, high inference latency, and difficulties of integration with large language models (LLMs) have necessitated exploring alternatives to DMs. This paper introduces ControlVAR, a novel framework that explores pixel-level controls in visual autoregressive (VAR) modeling for flexible and efficient conditional generation. In contrast to traditional conditional models that learn the conditional distribution, ControlVAR jointly models the distribution of image and pixel-level conditions during training and imposes conditional controls during testing. To enhance the joint modeling, we adopt the next-scale AR prediction paradigm and unify control and image representations. A teacher-forcing guidance strategy is proposed to further facilitate controllable generation with joint modeling. Extensive experiments demonstrate the superior efficacy and flexibility of ControlVAR across various conditional generation tasks against popular conditional DMs, e.g., ControlNet and T2I-Adaptor. Code: https://github.com/lxa9867/ControlVAR.
@@ 1 Introduction
In recent years, conditional image generation Zhang et al. (2023); Mou et al. (2023); Esser et al. (2021); Tian et al. (2024); Nam et al. (2024) has attracted great attention and there have been significant advancements in text-to-image generation Rombach et al. (2021); Chang et al. (2023); Gal et al. (2022), image-to-image generation Zhang et al. (2023); Mou et al. (2023); Ruiz et al. (2023), and even more complex tasks Nam et al. (2024); Li et al. (2023b; 2024). Most recent approaches, e.g., ControlNet Zhang et al. (2023), leverage the powerful diffusion models (DMs) Rombach et al. (2021); Peebles & Xie (2023) to model the large-scale image distribution and incorporate additional controls with classifier-free guidance Ho & Salimans (2022). However, the inherent nature of the diffusion process imposes many challenges for the diffusion-based visual generation: (1) the computational cost and inference time are significant due to the iterative diffusion steps Song et al. (2020a); Ho et al. (2020) and (2) the incorporation in mainstream intelligent systems, i.e., large language models (LLMs) Touvron et al. (2023); Achiam et al. (2023), is intricate due to the representation difference. This motivates the community to find a replacement for DMs for high-quality and efficient visual generation in the era of LLMs.
@@
Inspired by the success of autoregressive (AR) language modeling Touvron et al. (2023); Achiam et al. (2023), AR visual modeling Esser et al. (2021); Tian et al. (2024) has been investigated as a counterpart to DMs given its strong scalability and generalizability Tian et al. (2024); Bai et al. (2023). Several inspiring works, e.g., VQGAN Esser et al. (2021), DALL-E Ramesh et al. (2021a) and VAR Tian et al. (2024), have demonstrated promising image generation results with AR modeling. Nevertheless, compared to the prosperity of conditional DMs Zhang et al. (2023); Mou et al. (2023); Chen et al. (2022); Xu et al. (2023); Qin et al. (2023); Ju et al. (2023), visual generation with conditional AR modeling Zhan et al. (2022); Esser et al. (2021) remains significantly under-explored. Different from DMs, where all the pixels are modeled simultaneously, AR models are characterized by modeling sequential values based on their corresponding previous ones. This AR approach naturally leads to a conditional model, providing potential flexibility when incorporating additional controls. To leverage this property, teacher forcing is a popular approach that controls AR prediction by replacing partially predicted tokens with ground truth ones Esser et al. (2021). Thanks to this nature of AR modeling, we found that highly flexible conditional generation can be achieved by teacher forcing partial AR sequence with proper model designs.
@@
Figure 2: In contrast to previous methods Esser et al. (2021); Zhan et al. (2022) that leverage prefix conditional tokens to impose controls, ControlVAR jointly models the pixel-level controls and image during training and conducts the conditional generation tasks during testing with the teacher forcing. Class and type tokens provide semantic and control type (mask, canny, depth and normal) information respectively.
@@
In this paper, we explore the Controllable Visual AutoregRessive modeling with both token-level and pixel-level conditions. A new conditional AR paradigm, ControlVAR is introduced, which permits a highly flexible conditional image generation by embracing the next-scale prediction of joint control and image (Fig. 2(b)). Previous wisdom Zhan et al. (2022); Esser et al. (2021) typically utilizes prefix conditions (Fig. 2(a)) and mainly model images from raw pixel space in an AR manner. Differently, we notice that if we jointly model the control and image, the learned joint prediction can be easily guided by teacher forcing during inference. On the one hand, we unify the control and image representations and reformulate the sequential variables for the AR process to enable effective joint modeling. On the other hand, by analyzing the modeled probabilities, we introduce an effective sampling strategy, named teacher forcing guidance (TFG) to facilitate conditional sampling. Remarkably, a single ControlVAR model trained via TFG is capable of multiple meaningful tasks with different input-output combinations between control and image: (a) joint control-image generation, (b) control/image completion, (c) control-to-image generation, (d) image-to-control generation, as demonstrated in Fig. 1. Beyond the image-control tasks that are jointly modeled during training, we observe that ControlVAR also emerges capabilities for unseen tasks, e.g., control-to-control generation, further enhancing its flexibility and versatility. Our contribution can be summarized in three-fold:
@@ 2.1 Diffusion-based Image Generation
The evolution of diffusion models, initially introduced by Sohl-Dickstein et al. Sohl-Dickstein et al. (2015) and later expanded into image generation using fixed Gaussian noise diffusion processes Ho et al. (2020); Song et al. (2020b), has witnessed significant advancements driven by various research efforts. Nichol et al. Nichol & Dhariwal (2021) and Dhariwal et al. Dhariwal & Nichol (2021) proposed techniques to enhance the effectiveness and efficiency of diffusion models, paving the way for improved image generation capabilities. Notably, the paradigm shift towards modeling the diffusion process in the latent space of pre-trained image encoders as a strong prior Van Den Oord et al. (2017); Esser et al. (2021) rather than raw pixels spaces Vahdat et al. (2021); Rombach et al. (2022); Peebles & Xie (2023) has been instrumental in achieving high-quality image generation with reasonable inference speed. This approach has led to the development of foundational diffusion models such as Glide Nichol et al. (2021), Cogview Ding et al. (2021; 2022); Zheng et al. (2024), Make-a-scene Gafni et al. (2022b), Imagen Saharia et al. (2022), DALL.E Ramesh et al. (2021b), Stable Diffusion Stability AI (2022), MidJourney MidJourney Inc. (2022), SORA OpenAI (2024), among others, which are often pre-trained on large-scale data with conditions, typically text Gordon et al. (2023); Webster et al. (2023); Elazar et al. (2023); Chen et al. (2024). Recent advancements include consistency models derived from diffusion models Song et al. (2023); Song & Dhariwal (2023); Luo et al. (2023), enabling generation with reduced inference steps. These foundational diffusion models have not only opened doors to novel downstream applications like Text inversion Gal et al. (2022), DreamBooth Ruiz et al. (2023), T2I-Adapter Mou et al. (2023), ControlNet Zhang et al. (2023), but also inspired a plethora of research in controllable generation Meng et al. (2021); Brooks et al. (2023); Huang et al. (2023d); Tumanyan et al. (2023); Voynov et al. (2023); Huang et al. (2024; 2023a); Bashkirova et al. (2023); Bar-Tal et al. (2023); Li et al. (2023c); Qi et al. (2023); Zhan et al. (2022) and other innovative areas.
@@
Unlike diffusion-based models that typically leverage continuous image representation, autoregressive models Huang et al. (2023c); Esser et al. (2021); Van den Oord et al. (2016); Tian et al. (2024) leverage discrete image tokens. An image tokenizer Esser et al. (2021); Yu et al. (2023; 2024); Huang et al. (2023b); Ge et al. (2023) is utilized to encode the image into a sequence of discrete tokens. VQGAN Esser et al. (2021) first patches the image and then employs a vector-quantization approach to discretize the image features. Following this paradigm, a series of following-up works improve the image tokenization by using more powerful quantization operations Huang et al. (2023c); Lee et al. (2022); Yu et al. (2023), reformulating the image representation Tian et al. (2024); Tschannen et al. (2023) and modifying the network architecture Yu et al. (2021); Razavi et al. (2019). With the discrete tokens, a transformer structure Radford et al. (2019) is leveraged to model the image token sequences. RQ-GAN Lee et al. (2022) improves the modeling by incorporating a hierarchy design and MQ-VAE Huang et al. (2023c) further utilizes StackTransformer to enhance the spatial focus. MUSE Chang et al. (2023) is a large-scale pre-trained text-to-image model where a low-resolution image is first generated followed by a super-resolution transformer to refine the image. Recently, VAR Tian et al. (2024) introduced a new next-scale autoregressive prediction paradigm where the image representation is shifted from patch to scales. The new representation is featured with the maintenance of spatial locality and much lower computational cost. In this paper, we follow the next-scale autoregressive paradigm and explore the incorporation of additional controls into the modeling process.
@@ 2.3 Conditional Image Generation
Though significant progress has been made in generating highly realistic images from textual descriptions, describing every intricate detail of an image solely through text poses challenges. To overcome this limitation, researchers have explored alternative approaches using various additional inputs to effectively control image and video diffusion models. These inputs encompass bounding boxes Li et al. (2023c); Yang et al. (2023), reference object images Ruiz et al. (2023); Li et al. (2023a), segmentation maps Gafni et al. (2022a); Avrahami et al. (2023); Zhang et al. (2023), sketches Zhang et al. (2023), and combinations thereof Kim et al. (2023); Qin et al. (2023); Zhao et al. (2024); Wang et al. (2024); Mizrahi et al. (2024); Nam et al. (2024); Zhou et al. (2023). However, fine-tuning the vast array of parameters in these diffusion models can be computationally intensive. To address this, methods like ControlNet Zhang et al. (2023) have emerged, enabling conditional control through parameter-efficient training strategies Zhang et al. (2023); Ryu (2022); Mou et al. (2023). Notably, X-Adapter Ran et al. (2024) innovatively learns an adapter module to adapt ControlNets pre-trained on smaller image diffusion models (e.g., SDv1.5) for larger models (e.g., SDXL). SparseCtrl Guo et al. (2023) takes a different approach, guiding video diffusion models with sparse conditional inputs, such as few frames instead of full frames, to mitigate the data collection costs associated with video conditions. However, the implementation of SparseCtrl necessitates training a new variant of ControlNet from scratch, as it involves augmenting ControlNet with an additional channel for frame masks. Beyond traditional conditional image generation, the in-context learning capability of conditional models has also been explored Safaee et al. (2023); Mizrahi et al. (2024); Bai et al. (2023); Zhang et al. (2024). LVM Bai et al. (2023) investigates the scaling learning capability of a large vision model without any linguistic data. 4M Mizrahi et al. (2024) investigate the large-scale visual generation with multimodal data using masked image modeling. Different from previous works which are mainly focusing on diffusion models, we aim to explore adding additional control to the autoregressive visual generation process.
@@ 3 ControlVAR
ControlVAR is an autoregressive Transformer Vaswani et al. (2017) framework for conditional image generation tasks, using the following as conditions: image $I\in\mathbb{R}^{3\times H\times W}$, pixel-level control $C\in\mathbb{R}^{3\times H\times W}$ and token-level control $c\in\mathbb{R}^{D}$ where $H,W$ and $D$ denotes the image size and dimension of control token respectively. We denote the set of $N$ different types of controls as $\mathcal{C}=\{C_{n}\}_{n\in[N]}$.
@@
Prior conditional approaches Zhang et al. (2023); Tian et al. (2024) have often utilized distinct models for individual control type $C$, learning a conditional distribution in the form of $p(I|C,c)$, where each image $I$ is encoded as a sequence of discrete tokens of length $T$, denoted as $(x_{1},x_{2},\dots,x_{T})$. By employing autoregressive (AR) modeling, we can rewrite the conditional probability $p(I|C,c)$ as
@@
We consider four popular control types - entity mask, canny, depth, and normal in this paper. We notice that canny, depth, and normal can be easily converted to RGB by using simple transformations Zhang et al. (2023). However, entity segmentation masks $M\in\{0,1\}^{N\times H\times W}$ which comprises $N$ class-agnostic binary masks (Fig. 3(b)) cannot be easily converted.
@@
Inspired by SOLO Wang et al. (2020), we leverage a position-aware color map to encode the binary masks $M$ into a colormap $M^{\prime}\in[0,255]^{3\times H\times W}$. To better distinguish the color difference, we select 5 candidate values $\{0,64,128,192,255\}$ from each RGB channel and combine them to $1124=5^{3}-$ 1 colors ($(0,0,0)$ is preserved for background). To apply the colormap, as shown in Fig. 3, we divide the image into $n_{h}\times n_{w}$ regions where each region represents a corresponding color. We calculate the centeredness of each mask and apply the colors to masks based on their centeredness locations. Therein, we can convert the entity masks to a RGB colormap.
[…]
@@
As the control and image share the same RGB representation, we can utilize the same approach to tokenize them. To represent an RGB image as a sequence of discrete tokens $(x_{1},x_{2},\dots,x_{T})$, patch-level Esser et al. (2021) and scale-level Tian et al. (2024) representations have been explored. The patch-level tokenization process splits an image into $T$ patches and represents each patch as a token $x_{t}$ where $x_{t}\in[V]^{1}$ is an integer from a vocabulary of size $V$. Recently, a scale-level representation has been introduced which decomposes the image into $T$ scales where each scale is represented by a set of tokens $x_{t}\in[V]^{h_{t}\times w_{t}}$ (Fig. 2(b)). $h_{t}\times w_{t}$ denotes the size of the $t$-th scale. Compared to patch-level representation, scale-level representation can better preserve the spatial locality and capture global information which are desired for conditional image generation tasks. This motivates us to adopt the scale-level representation in our approach. Specifically, we obtain the image tokens and control tokens using the shared tokenizer $\Phi$ as
@@ 3.2 Joint Control-Image Modeling
We demonstrate the network details for joint modeling in this section. Following VAR Tian et al. (2024), we leverage a GPT-2 style Transformer network architecture for our ControlVAR models. As shown in Fig. 4 (a), we jointly model the control and image in each stage. A flatten operation is adopted to convert the sequence of 2D features into 1D. Full attention is enabled for both control and image tokens belonging to the same scale, which allows the model to maintain spatial locality and to exploit the global context between control and image. A standard cross entropy loss is used to supervise our autoregressive ControlVAR models.
Specifically, we employ two pre-defined special tokens $c=\texttt{[CLS]}\in[N_{cls}]^{1}$ and $c_{t}=\texttt{[TYP]}\in[N_{typ}]^{1}$ as the start tokens. $N_{cls}$ and $N_{tpy}$ denote the number of classes and control types respectively. [CLS] token aims to provide semantic context for the generated image. [TYP] token is used to select the type of pixel-level control to be generated along with the image. Following previous works Chang et al. (2023); Tian et al. (2024), additional empty tokens are used to replace special tokens with a probability of $\delta$ during training to apply classifier-free guidance Ho & Salimans (2022).
@@
Classifier-free guidance (CFG) Ho & Salimans (2022) was originally introduced to apply and enhance the effect of conditional controls on diffusion models without an explicit classifier. Extensive studies Sanchez et al. (2023); Chang et al. (2023); Tian et al. (2024) have demonstrated that classifier-free guidance also works for AR models.
@@
Here, we analyze how to achieve conditional generation by using the image generation task $p(I|C,c,c_{t})$ as an example. Given image $I$, pixel-level control $C$ and token-level controls $c,c_{t}$, CFG Ho & Salimans (2022) leverages Bayesian rule to rewrite the conditional distribution as